OpenStudy (anonymous):

hi, i'm integrating (2x^2 +3)/(x^2 +1)^2 and i'd reach the following step and get stucked ==' 3/2tan^-1 (x) + 1/4 sin [2(tan^-1 (x))]

OpenStudy (anonymous):

basically its the sin of 2u, the u i previously let to be tan (x), now it became sin of 2 tan^-1 x

OpenStudy (anonymous):

my calculus foundation sux to the max ==

OpenStudy (anonymous):

What I would do is break it up: \[2 \int\limits \frac{x^2 dx}{(x^2+1)^2}+3 \int\limits \frac{dx}{(x^2+1)^2}\] From here you can do trig sub: Let \[\tan(\Omega)=x; \sec^2(\Omega) d \Omega= dx; \sec(\Omega)=\sqrt{x^2+1}\] All of that is coming from the pythagorean theorem. Replace this in the integral \[\sec(\Omega)=\sqrt{x^2+1} \implies \sec^4(\Omega)=(x^2+1)^2\] \[2 \int\limits \frac{\tan^2(\Omega)\sec^2(\Omega)}{\sec^4(\Omega)}d \Omega+3 \int\limits \frac{\sec^2(\Omega)}{\sec^4(\Omega)} d \Omega\] \[2 \int\limits \frac{\sin^2(\Omega)}{\cos^2(\Omega)}\cos^2(\Omega) d \Omega + 3 \int\limits \cos^2(\Omega) d \Omega\] Do you follow any of this?

OpenStudy (anonymous):

I guess partial fractions would have been easier >.>

OpenStudy (anonymous):

yea i approached it with partial fractions, the integral of the 1st term with x^2 +1 i got 2 tan^-1 x, but i think im stucked at the (x^2+1)^2 part

OpenStudy (anonymous):

erm, how the 3 is formed at the 2nd partial fraction term? i got 1 on top o.o

OpenStudy (anonymous):

Okay, the do partial fractions: \[\frac{2x^2+3}{(x^2+1)^2}=\frac{Ax+\eta}{x^2+1}+\frac{\xi x + B}{(x^2+1)^2}\] Getting a common denominator implies that \[A=0; B=0; \eta = 2; \xi = 1\] Gives: \[I=2 \int\limits \frac{dx}{x^2+1}+ \int \frac{dx}{(x^2+1)^2}\] The first term is indeed 2 arctan: \[I=2 \arctan(x)+\int\limits \frac{dx}{(x^2+1)^2}\]

OpenStudy (anonymous):

Is this where you are stuck?

OpenStudy (anonymous):

not sure whether the 2nd part i integrated wrongly or not, cant merge all the 3 terms 2nd part i get \[1/4 \sin 2\left( \tan^{-1} x \right) + 1/2 \tan^{-1} \]

OpenStudy (anonymous):

or its just me, dunoe how to simplify them ==

OpenStudy (anonymous):

 Also: tan(omega)=opp/adj=x So: \[\sec^2(\Omega) \frac{d \Omega}{dx}=1 \implies \sec^2(\Omega) d \Omega= dx\] Also sec(omega)=1/cos(omega)=1/(adj/hyp)=hyp/adj=sqrt{1+x^2}/1 So we have: \[\tan(\xi)=x; \sec^2(\xi) d \xi = dx; \sec(\xi)=\sqrt{1+x^2}\]